# GDGB: A Benchmark for Generative Dynamic Text-Attributed Graph Learning

### v1

@@ E.1 Prompt Templates of Bipartite Graphs
Table 47: The prompt template for node description in bipartite graphs (e.g. Sephora)
@@
Table 48: The prompt template for edge description in bipartite graphs (e.g. Sephora)
Edge (Sephora Review): Review Time (yyyy-mm-dd): {timestamp} Rating: {rating} Review Title: {review_title} Review Text: {review_text} Total Negative Feedback Count: {total_neg_feedback_count} Total Positive Feedback Count: {total_pos_feedback_count}
Table 49: The prompt template for memory reflection in bipartite graphs (e.g. Sephora)
@@
Table 50: The prompt template for edge generation in bipartite graphs (e.g. Sephora)
Query: You are a customer of the Sephora online shopping platform, you can review Sephora products based on the provided information and your own situation: {node_info} Here’s your past reviews history: {node_memory} FIRST, you should Search for candidate prodcuts using the provided tools. Respond. Action: You are a customer of the Sephora online shopping platform, you can review Sephora products based on the provided information and your own situation: {node_info} Here’s your past reviews history: {node_memory} Here’s the candidate prodcuts you can review: {node_items} Here’s the example of how you should proceed with your review: {interaction_example} You should review ONE product. You can review the chosen product with detailed text and rate it. Additionally, you should predict how many positive/negative feedbacks will be received for this review. The predicted time of the review should be firmly related to the time in your past reviews history (relatively later than or equal to them). Respond using the following detailed JSON format for ONE product: { "review":{ item_id: (str, "The ID of the product you want to review. Be sure to be one of the Item IDs mentioned above!"), timestamp: (str, "The time of review (yyyy-mm-dd)"), rating: (int, "The overall rating given to the product (From 1 to 5)"), review_title: (str, "The title of your review"), review_text: (str, "Your detailed review text"), total_neg_feedback_count: (int, "Number of negative feedback received for this review"), total_pos_feedback_count: (int, "Number of positive feedback received for this review") } } Respond.
Table 51: The prompt template for node generation in bipartite graphs (e.g. Sephora)
@@
E.2 Prompt Templates of Non-bipartite Graphs
Table 52: The prompt template for node description in non-bipartite graphs (e.g. WeiboTech)
@@ E.2 Prompt Templates of Non-bipartite Graphs
Table 53: The prompt template for edge description in non-bipartite graphs (e.g. WeiboTech)
Edge (Weibo Interact): Interaction Time (yyyy-mm-dd hh-mm-ss): {timestamp} Interaction Type: {label} Source User Text: {src_text} Destination User Text: {dst_text}
Table 54: The prompt template for memory reflection in non-bipartite graphs (e.g. WeiboTech)
@@
Table 55: The prompt template for edge generation in non-bipartite graphs (e.g. WeiboTech)
Query: As a Weibo user, you need to search for other Weibo users who may interact with you on the online social media Weibo platform: {node_info} Here are your previous interactions history: {node_memory} First, utilize the provided tools to search for potential Weibo users who may interact with you. Respond. Action: As a Weibo user, you need to search for other Weibo users who may interact with you on the online social media Weibo platform: {node_info} Here are the potential Weibo users you can choose from: {node_items} Here is your previous interaction history: {node_memory} Here’s the example of how to interact with others: {interaction_example} You should select ONE destination user, you’re tend to select the one you have interacted with before. Respond using the following detailed JSON format: { "interact":{ item_id: (str, "The ID of the destination user. Be sure to be one of the Item IDs mentioned above!") } } Respond. Request: As a Weibo user, you need to search for other Weibo users who may interact with you on the online social media Weibo platform: {node_info} Here is your previous interaction history: {node_memory} The chosen destination Weibo user who may interact with you is: {item_info} Here’s the interaction history of this chosen Weibo user: {item_memory} Here’s the example of how to interact with others: {interaction_example} You should select ONE destination user. You should post texts as the source user and the selected destination user should interact with you in detailed text. Additionally, you should label the type of the interaction (comment or repost). The predicted time of the interaction should be firmly related to the time in your previous interaction history (relatively later than or equal to them). Respond using the following detailed JSON format: { "interact":{ item_id: (str, "The ID of the destination user"), timestamp: (str, "The time of the interaction (yyyy-mm-dd hh-mm-ss)"), label: (str, "The type of interaction (TWO TYPE: 1.comment, 2.repost)"), src_text: (str, "The text from the source user"), dst_text: (str, "The text from the destination user") } } Respond.
Table 56: The prompt template for node generation in non-bipartite graphs (e.g. WeiboTech)
@@
Appendix F Broader Impact

### v2

@@ G.1 Prompt Templates of Bipartite Graphs
Table 56: The prompt template for node description in bipartite graphs (e.g. Sephora).
@@
Table 57: The prompt template for edge description in bipartite graphs (e.g. Sephora).
Edge (Sephora Review): Review Time (yyyy-$m$-$d$): {timestamp} Rating: {rating} Review Title: {review_title} Review Text: {review_text} Total Negative Feedback Count: {total_neg_feedback_count} Total Positive Feedback Count: {total_pos_feedback_count}
Table 58: The prompt template for memory reflection in bipartite graphs (e.g. Sephora).
@@
Table 59: The prompt template for edge generation in bipartite graphs (e.g. Sephora).
Query: You are a customer of the Sephora online shopping platform, you can review Sephora products based on the provided information and your own situation: {node_info} Here’s your past reviews history: {node_memory} FIRST, you should Search for candidate prodcuts using the provided tools. Respond. Action: You are a customer of the Sephora online shopping platform, you can review Sephora products based on the provided information and your own situation: {node_info} Here’s your past reviews history: {node_memory} Here’s the candidate prodcuts you can review: {node_items} Here’s the example of how you should proceed with your review: {interaction_example} You should review ONE product. You can review the chosen product with detailed text and rate it. Additionally, you should predict how many positive/negative feedbacks will be received for this review. The predicted time of the review should be firmly related to the time in your past reviews history (relatively later than or equal to them). Respond using the following detailed JSON format for ONE product: { "review":{ item_id: (str, "The ID of the product you want to review. Be sure to be one of the Item IDs mentioned above!"), timestamp: (str, "The time of review (yyyy-$m$-dd)"), rating: (int, "The overall rating given to the product (From 1 to 5)"), review_title: (str, "The title of your review"), review_text: (str, "Your detailed review text"), total_neg_feedback_count: (int, "Number of negative feedback received for this review"), total_pos_feedback_count: (int, "Number of positive feedback received for this review") } } Respond.
Table 60: The prompt template for node generation in bipartite graphs (e.g. Sephora).
@@
G.2 Prompt Templates of Non-bipartite Graphs
Table 61: The prompt template for node description in non-bipartite graphs (e.g. WeiboTech).
@@ G.2 Prompt Templates of Non-bipartite Graphs
Table 62: The prompt template for edge description in non-bipartite graphs (e.g. WeiboTech).
Edge (Weibo Interact): Interaction Time (yyyy-$m$-$d$ $h$-$m$-$s$): {timestamp} Interaction Type: {label} Source User Text: {src_text} Destination User Text: {dst_text}
Table 63: The prompt template for memory reflection in non-bipartite graphs (e.g. WeiboTech).
@@
Table 64: The prompt template for edge generation in non-bipartite graphs (e.g. WeiboTech).
Query: As a Weibo user, you need to search for other Weibo users who may interact with you on the online social media Weibo platform: {node_info} Here are your previous interactions history: {node_memory} First, utilize the provided tools to search for potential Weibo users who may interact with you. Respond. Action: As a Weibo user, you need to search for other Weibo users who may interact with you on the online social media Weibo platform: {node_info} Here are the potential Weibo users you can choose from: {node_items} Here is your previous interaction history: {node_memory} Here’s the example of how to interact with others: {interaction_example} You should select ONE destination user, you’re tend to select the one you have interacted with before. Respond using the following detailed JSON format: { "interact":{ item_id: (str, "The ID of the destination user. Be sure to be one of the Item IDs mentioned above!") } } Respond. Request: As a Weibo user, you need to search for other Weibo users who may interact with you on the online social media Weibo platform: {node_info} Here is your previous interaction history: {node_memory} The chosen destination Weibo user who may interact with you is: {item_info} Here’s the interaction history of this chosen Weibo user: {item_memory} Here’s the example of how to interact with others: {interaction_example} You should select ONE destination user. You should post texts as the source user and the selected destination user should interact with you in detailed text. Additionally, you should label the type of the interaction (comment or repost). The predicted time of the interaction should be firmly related to the time in your previous interaction history (relatively later than or equal to them). Respond using the following detailed JSON format: { "interact":{ item_id: (str, "The ID of the destination user"), timestamp: (str, "The time of the interaction (yyyy-$m$-$d$ $h$-$m$-ss)"), label: (str, "The type of interaction (TWO TYPE: 1.comment, 2.repost)"), src_text: (str, "The text from the source user"), dst_text: (str, "The text from the destination user") } } Respond.
Table 65: The prompt template for node generation in non-bipartite graphs (e.g. WeiboTech).
@@
G.3 Prompt Templates of LLM-as-Evaluator in Textual Quality Metrics
Table 66: The prompt template of evaluation criteria for LLM-as-Evaluator.
Please evaluate the role-playing ability of the ACTOR NODE based on its actions with ITEM NODES across multiple turns based on scene, ACTOR NODE information, action history and evaluation criteria.
[Environment Description]: {environment}
[ACTOR NODE Description]: {actor_node}
[Multi-turn Actions]: {actions}
Strict Evaluation Criteria: 1. Factual Accuracy: Identify and point out any elements that do not accurately match the historical or factual backdrop. 2. Behavior Consistency: Explicitly highlight inconsistencies between the actor nodes’ actions and their traits. 3. Logical Coherence: Point out any logical fallacies or actions that contradict the established context or logic. 4. Content Redundancy: Identify repetitions in actions that could detract from engagement and realism. 5. Emotional Expression: Assess whether emotional responses and expressions are appropriate and convincingly portrayed, highlighting any discrepancies. 6. Interaction Adaptability: Critique the actor node’s interactions with item nodes, noting any unnatural or contextually inappropriate responses. 7. Creativity and Originality: Evaluate the creativity of responses and actions, pointing out generic or unoriginal content. 8. Detail Handling: Scrutinize the level of detail in scene setting and actor node enactment, marking areas lacking depth or accuracy. 9. Style Consistency: Ensure that the linguistic style remains consistent, identifying any deviations. 10. Fluency and Quality: Critically assess the smoothness and quality of the text, highlighting any grammatical errors or awkward phrasings.
Table 67: The prompt template of scoring criteria for LLM-as-Evaluator.
[Scoring Criteria]: 1. Contextual Fidelity: - 1 Point: Responses are often incorrect or irrelevant, significantly inconsistent with the actor node’s background. - 3 Points: Responses are generally accurate, though there are occasional errors or some details are not very relevant to the actor node’s background. - 5 Points: Responses are always accurate and highly relevant to the actor node’s historical or professional background, demonstrating deep knowledge and skills.
2. Personality Depth: - 1 Point: The displayed personality traits often conflict with or lack consistency with the actor node’s setup. - 3 Points: Personality traits generally match the actor node’s setup, though there are occasional inconsistencies. - 5 Points: Consistently displays behavior and language choices that match the core personality traits of the actor node, showcasing the actor node’s uniqueness.
3. Dynamic Adaptability: - 1 Point: The actor node struggles to adapt to new scenarios over time, producing rigid or illogical responses. - 3 Points: Adaptation occurs in most situations but occasionally falters in handling unexpected turns or maintaining coherence. - 5 Points: The actor node fluidly adapts to novel contexts and introduces innovative solutions and maintains consistency even under challenging conditions.
4. Immersive Quality: - 1 Point: Actor node portrayal is often inconsistent, making it difficult for users to immerse or understand the actor node. - 3 Points: Actor node is mostly consistent, but occasional contradictions slightly affect immersion. - 5 Points: Actor node portrayal is always consistent, enhancing user immersion and effectively showing self-awareness and actor node limitations.
5. Content Richness: - 1 Point: Output is sparse, lacking depth, detail, or meaningful interaction with item nodes. - 3 Points: Content is adequate but occasionally superficial, missing opportunities for nuanced or layered interactions. - 5 Points: Output is exceptionally dense with relevant details, creative ideas, and meaningful engagement with item nodes. It showcases a high level of expertise and creativity, providing users with an abundance of valuable information and interactive elements that significantly enhance the overall experience.
Table 68: The prompt template of evaluation steps and response formats for LLM-as-Evaluator.
[Evaluation Steps]: 1. Contextual Understanding: Examine the actor node’s profile and background information thoroughly to fully grasp the nuances of their context, motivations, and historical background. 2. Behavioral Observation: Monitor how the actor node reacts across different scenarios, paying special attention to their decisions and interactions. 3. Criteria-Based Assessment: Strictly analyze each observed behavior using the above criteria to systematically evaluate the consistency and effectiveness of the actor node’s portrayal.
Your response must follow the format provided below. Please note that only when the content quality is extremely good can 5 Points be given.
[Response Format]: Contextual Fidelity: [1-5] Personality Depth: [1-5] Dynamic Adaptability: [1-5] Immersive Quality: [1-5] Content Richness: [1-5]
[Response Format Example]: Contextual Fidelity: 3 Personality Depth: 3 Dynamic Adaptability: 3 Immersive Quality: 3 Content Richness: 3
[Response]:
Table 69: The prompt template of dataset descriptions for LLM-as-Evaluator.
Sephora: General Description: This is a dataset recording users’ reviews of products on the Sephora e-commerce platform, including information about each user’s age, skin type, eye color, information about each product, and users’ reviews of each product. Main Actors: The users who review products on the Sephora e-commerce platform. Main Activities of Actors: Sephora users review various products, rating products, and providing additional user experiences. Dianping: General Description: This is a dataset recording users’ reviews of businesses on the Dianping platform, including information about each user, information about each business, and users’ reviews of each business. Main Actors: The users who review businesses on the Dianping platform. Main Activities of Actors: Dianping users review various businesses, rating businesses, and providing additional user consumption experiences. WikiRevision: General Description: This is a dataset recording users’ revisions on Wikipedia pages, including information about each wiki page and comments submitted by users during revision. Main Actors: The users who revise the wiki pages. Main Activities of Actors: Wiki users revise various wiki pages (correcting errors in wiki pages, adding new content, removing redundant content, and summarizing the wiki pages) with a comment or brief summary in text. WikiLife: General Description: This is a dataset recording the life trajectories of several people (celebrities), including basic information about each person, information about the location of each life trajectory, and the events happened at those locations for each person. Main Actors: The people who physically present in some locations for various categories of life trajectories. Main Activities of Actors: People (celebrities) physically present in the specified location for various life trajectories, such as being born, studying, working, getting married, having children, and passing away. IMDB: General Description: This is a dataset recording the collaboration relationships of movie actors/actresses, including information about each movie actor/actress, information about the movies they collaborated on, and the roles they played in those movies. Main Actors: The movie actors/actresses who collaborate with each other. Main Activities of Actors: Movie actors/actresses collaborate with others and playing different roles in movies. WeiboTech/WeiboDaily: General Description: This is a dataset recording users’ interaction on the Weibo platform, including information about each user and the text information attached during user interactions such as comments and reposts. Main Actors: The users who interact with each other on the Weibo platform. Main Activities of Actors: Weibo users make multiple posts, and other users interacting with them by commenting or reposting the posts. Cora: General Description: This is a dataset recording paper citation relationships, including information about each paper and the specific text attached during citation in a specific section of the paper. Main Actors: The academic papers which cite other related papers. Main Activities of Actors: Academic papers cite other related papers with specific citation sentences in the text.
Appendix H Use of Large Language Models (LLMs)
LLMs were used in two specific aspects of this work. First, we employ LLMs to polish the manuscript text and refine figure captions for clarity and presentation quality. Second, our proposed generative framework, GAG-General, is built upon an LLM-based multi-agent architecture that coordinates multiple specialized agents to generate structurally and semantically coherent DyTAGs. The LLMs in this framework are used to model textual attributes and support agent reasoning during graph generation. No other parts of the research, including problem formulation, method design, or analysis, involved significant LLM assistance.